OpenStudy (anonymous):

3/6x +3/8 =3/8x +2/6

OpenStudy (anonymous):

x = -3

OpenStudy (anonymous):

First simplify the fractions (see the attachment) then separate the x terms and the numbers

OpenStudy (anonymous):

i think i told you wrong its (3/6)x +3/8 =(3/8)x +2/6

OpenStudy (anonymous):

ok then x = -1/3

OpenStudy (anonymous):

thats what i thought but it says its incorrect

OpenStudy (anonymous):

:S should be right though